# Improving Implicit Regularization of SGD with Preconditioning for Least Square Problems

### v1

@@ Abstract
Stochastic gradient descent (SGD) exhibits strong algorithmic regularization effects in practice and plays an important role in the generalization of modern machine learning. However, prior research has revealed instances where the generalization performance of SGD is worse than ridge regression due to uneven optimization along different dimensions. Preconditioning offers a natural solution to this issue by rebalancing optimization across different directions. Yet, the extent to which preconditioning can enhance the generalization performance of SGD and whether it can bridge the existing gap with ridge regression remains uncertain. In this paper, we study the generalization performance of SGD with preconditioning for the least squared problem. We make a comprehensive comparison between preconditioned SGD and (standard & preconditioned) ridge regression. Our study makes several key contributions toward understanding and improving SGD with preconditioning. First, we establish excess risk bounds (generalization performance) for preconditioned SGD and ridge regression under an arbitrary preconditions matrix. Second, leveraging the excessive risk characterization of preconditioned SGD and ridge regression, we show that (through construction) there exists a simple preconditioned matrix that can outperform (standard & preconditioned) ridge regression. Finally, we show that our proposed preconditioning matrix is straightforward enough to allow robust estimation from finite samples while maintaining a theoretical advantage over ridge regression. Our empirical results align with our theoretical findings, collectively showcasing the enhanced regularization effect of preconditioned SGD.
@@ 1 Introduction
A growing body of work has emerged, dedicated to understanding and characterizing the generalization performance of SGD in the context of the linear regression problem. These studies have unveiled a close relationship between the implicit regularization inherent to SGD and the explicit $L_{2}$-type regularization (ridge) (Zhang et al., 2017; Tikhonov, 1963; Ali et al., 2019; 2020; Gunasekar et al., 2018; Suggala et al., 2018; Zou et al., 2021a). In particular,  Zou et al. (2021a) has presented an instance-based comparison study between the generalization performance (measured by excessive risk) of SGD and ridge regression. They show that (single-pass) SGD can achieve better generalization performance than that of ridge regression in certain problem instances but fall behind in others.
A notable insight derived from the comparative analysis conducted by Zou et al. (2021a) is that the underperformance of SGD in certain scenarios can be attributed to an inherent imbalance optimization within the data covariance matrix. This imbalance results in a substantial bias error, primarily associated with the “relatively small” eigenvalues of the data covariance matrix. To address this limitation of SGD, preconditioning emerges as a natural solution. Preconditioning is a commonly used technique to enhance the performance of the optimizer (Benzi, 2002; Avron et al., 2017; Gupta et al., 2018; Amari et al., 2020; Kang et al., 2023; Li, 2017; Li et al., 2016); it involves modifying the update steps of the algorithm by transforming the data space, offering a mechanism to rebalance the optimization process across different directions. Despite extensive research that has explored the implicit regularization of SGD and its parallels with ridge regression, the realm of preconditioned SGD remains relatively uncharted. Recognizing this gap in the current body of knowledge, our study investigates the implicit regularization of SGD when augmented with preconditioning. We adhere to the standard settings used in the literature (least square problem) (Zou et al., 2021a; b; 2023; Tsigler & Bartlett, 2023) and aim to address the following central question of our study:
@@
In essence, our objective is to investigate whether a thoughtfully designed preconditioning matrix can consistently empower SGD to outperform ridge regression in the least square problem. Furthermore, we aim to delve deeper into the dynamic relationship between SGD and ridge regression when both methods incorporate preconditioning. Achieving these goals presents several key challenges. First and foremost, existing results characterizing the excessive risk (generalization performance) of SGD and ridge regression predominantly pertain to scenarios without preconditioning. The introduction of preconditioning alters the learning dynamics of SGD and ridge regression, potentially leading to distinct characterizations of their excessive risk. Secondly, as discussed earlier, the effectiveness of preconditioning for SGD is contingent upon its impact on the eigenspectrum of the data covariance matrix. A naive design of the preconditioning matrix may fail to align optimally with the excessive risk characteristics of SGD, potentially leading to adverse consequences. Lastly, obtaining precise information regarding the eigenspectrum of the covariance matrix can be impractical in real-world settings. To ensure the practical applicability and validity of the preconditioning technique in such scenarios, we must consider a design that remains sufficiently straightforward to facilitate robust estimation.
@@
In this paper, we show (through construction) that there exists a preconditioning design that allows the excessive risk of SGD to consistently outperform ridge regression. In addition, we establish corresponding analytical tools to tackle the challenges stated earlier. The main contributions of our work are summarized as follows.
@@
  • Building upon the insights derived from the analysis of excessive risk, we propose a simple (yet effective) preconditioning matrix tailored to SGD, which leverages information from the data covariance matrix. We show that under the theoretical setting (with access to the precise information of the data covariance matrix), the excessive risk of SGD with our proposed preconditioning can consistently outperform the standard ridge regression and a representative family of preconditioned ridge regression (Theorems 4.3 &  4.4). This shows that preconditioning can indeed bridge the remaining gap between SGD and ridge regression, supporting the assertion of the central question above. In addition, the results suggest that the advantages of preconditioning can be harnessed to a greater extent by SGD.
  • In practical scenarios where precise information about the data covariance matrix is unavailable, our study illustrates that our proposed preconditioning design allows for robust estimation using a finite set of inexpensive unlabeled data. Remarkably, even when we employ the estimated version of our proposed preconditioning matrix, SGD can still consistently outperform the excessive risk of (standard & preconditioned) ridge regression (Theorems 4.5 &  4.6). This discovery not only highlights the significance of SGD in practical machine learning applications but also underscores the practical advantages offered by preconditioning.
Notation. We use lowercase letters to denote scalars, and use lower and uppercase boldface letters to denote vectors and matrices, respectively. For a vector $\mathbf{x}$, $\mathbf{x}[i]$ denotes the $i$-th coordinate of $\mathbf{x}$. For two functions $f(x)\geq 0$ and $g(x)\geq 0$ defined on $x>0$, we write $f(x)\lesssim g(x)$ if $f(x)\leq c\cdot g(x)$ for some absolute constant $c>0$; we write $f(x)\gtrsim g(x)$ if $g(x)\lesssim f(x)$; we write $f(x)\eqsim g(x)$ if $f(x)\lesssim g(x)\lesssim f(x)$. For a vector $\mathbf{w}\in\mathbb{R}^{d}$ and a positive semidefinite matrix $\mathbf{H}\in\mathbb{R}^{d\times d}$, we denote $\|\mathbf{w}\|^{2}_{\mathbf{H}}:=\mathbf{w}^{\top}\mathbf{H}\mathbf{w}$.
@@ 2 Related Work
In this section, we provide an overview of the technical advancements in the analysis of ridge regression and SGD, as well as related research comparing SGD to explicit norm-based regularization and ridge regression. Additionally, we discuss the current usage of preconditioning in both SGD and ridge regression.
2.1 Excessive Risk Bounds of Ridge Regression
@@
Ridge regression holds a central position in the field of machine learning (Hsu et al., 2012; Kobak et al., 2020; Tsigler & Bartlett, 2023; Hastie et al., 2009). In the underparameterized scenario, there exists comprehensive understanding on the excess risk bounds of ridge regression (Hsu et al., 2012). However, in the overparameterized regime, a large body of work focuses on characterizing the excess risk of ridge regression in the asymptotic scenario where both the sample size and the dimension approach infinity with a constant ratio (Dobriban & Wager, 2018; Hastie et al., 2022; Wu & Xu, 2020; Xu & Hsu, 2019). Recent advancements by (Tsigler & Bartlett, 2023) have provided detailed non-asymptotic risk bounds for ordinary least squares and ridge regression in overparameterized contexts. Notably, these bounds depend on the data covariance spectrum and are applicable even when the ridge parameter is zero or negative. This body of work forms a foundation of the theoretical framework for our paper.
2.2 Excessive Risk Bounds of SGD
@@
In the finite-dimensional setting, risk bounds for one-pass, constant-stepsize SGD have been well-established (Dieuleveut et al., 2017; Bach & Moulines, 2013; Jain et al., 2017; 2018; Défossez & Bach, 2015). In the under-parameterized setting, constant-stepsize SGD with tail-averaging has been recognized for achieving the minimax optimal rate for least squares (Jain et al., 2017; 2018). A recent study by (Zou et al., 2023) expanded on the previous analysis, offering a detailed risk bound for SGD in the overparameterized regime. This work provides closely aligned upper and lower excess risk bounds for constant-stepsize SGD, distinctly defined in relation to the complete eigenspectrum of the population covariance matrix. These findings are critically influential in the context of our paper.
2.3 Implicit Regularization of SGD and Explicit Regularization
@@
It is well known that in the least squares problems, multi-pass SGD approaches the solution with the minimum norm, representing a recognized implicit bias of SGD (Neyshabur et al., 2014; Zhang et al., 2017; Gunasekar et al., 2018). However, this norm-based regularization alone does not fully explain the optimization behaviour of SGD in more extensive scenarios such as convex but non-linear models (Arora et al., 2019; Dauber et al., 2020; Razin & Cohen, 2020). Recent work by (Zou et al., 2021a) presents a comparative analysis between SGD and ridge regression from the perspective of excessive risk, demonstrating that SGD can outperform ridge regression in some scenarios but fall short in others. Their result has motivated us to explore SGD with preconditioning. In this paper, we advance the understanding of SGD by exploring its implicit regularization in the presence of preconditioning.
2.4 Preconditioning in SGD and ridge
@@
Preconditioning is a widely employed optimization technique (Benzi, 2002) aimed at enhancing the efficiency and effectiveness of optimization algorithms. Its primary objective is to transform complex or ill-conditioned optimization problems into more manageable forms. Preconditioning techniques have been extensively explored in both SGD and ridge regression, serving various purposes (Benzi, 2002; Li, 2017; Gonen et al., 2016; Amari et al., 2020; Xu et al., 2023; Kang et al., 2023). However, the impact of preconditioning on the excessive risk of SGD and ridge regression, and whether a preconditioning matrix can enable SGD to consistently outperform ridge regression, remain open questions that motivate our research.
@@ 3 Problem Setup and Preliminaries
In this paper, our objective is to explore the usage of precondition in SGD and compare the generalization performance of preconditioned SGD and ridge regression in the context of solving least square problems in the overparameterized regime. We denote a feature vector in a separable Hilbert space $\mathcal{H}$ as $\mathbf{x}\in\mathcal{H}$. The dimensionality of $\mathcal{H}$ is denoted as $d$, which can be infinite-dimensional when applicable. We use $y\in\mathbb{R}$ to denote a response and it is generated as follows:
@@
where $\mathbf{w}^{*}\in\mathcal{H}$ represents the unknown true model parameter, and $\epsilon\in\mathbb{R}$ is the model noise. The goal of the least squares problem is to estimate the true parameter $\mathbf{w}^{*}$ by minimizing the following objective function:
$L(\mathbf{w}^{*})=\min_{\mathbf{w}\in\mathcal{H}}L(\mathbf{w}),$
$L(\mathbf{w}):=\frac{1}{2}\mathbb{E}_{(\mathbf{x},y)\sim\mathbb{D}}\big{[}(y-\langle\mathbf{w},\mathbf{x}\rangle)^{2}\big{]},$
where $\mathbb{D}$ represents the data distribution. The generalization performance of an estimation $\mathbf{w}$ found by an algorithm, whether it is SGD or ridge regression, is evaluated based on the excessive risk, given by:
@@
There exists a positive constant $\alpha$ such that for any PSD matrix $\mathbf{A}$, it holds that:
$\mathbb{E}[\mathbf{x}\mathbf{x}^{\top}\mathbf{A}\mathbf{x}\mathbf{x}^{\top}]-\mathbf{H}\mathbf{A}\mathbf{H}\preceq\alpha\mathop{\mathrm{tr}}(\mathbf{A}\mathbf{H})\cdot\mathbf{H}.$
@@
There exists a positive constant $\sigma$ such that:
$\mathbb{E}_{\mathbf{x},\epsilon}[\epsilon^{2}\mathbf{x}\mathbf{x}^{\top}]\leq\sigma^{2}\mathbf{H}.$
To analyze the relationship between $\mathbf{w}^{*}$ and $\mathbf{H}$ in terms of excessive risk, we introduce the following notation:
@@
Preconditioned SGD. We consider single-pass preconditioned SGD with a constant step size and tail-averaging (Bach & Moulines, 2013; Jain et al., 2017; 2018). At the $t$-th iteration, a fresh example $(\mathbf{x}_{t},y_{t})$ is sampled from the data distribution and to perform the update:
${\mathbf{w}}_{t+1}$ $=\mathbf{w}_{t}-\eta\cdot{\color[rgb]{0,0,1}{\mathbf{G}}}\nabla l(\mathbf{w}_{t};\mathbf{x}_{t},y_{t})$
$={\mathbf{w}}_{t}-\eta\cdot(\langle{\mathbf{w}}_{t},\mathbf{x}_{t}\rangle-y_{t})\cdot{\color[rgb]{0,0,1}{\mathbf{G}}}\mathbf{x}_{t}$ (3.1)
where $\eta>0$ is a constant stepsize (also referred to as the learning rate) and $\mathbf{G}$ (highlighted in blue) is the preconditioned matrix — the key difference from the standard SGD. Note that if $\mathbf{G}=\mathbf{I}$, then the update rule (3) reduces to the standard SGD with constant step size.
@@
$\mathbf{w}_{\mathrm{sgd}}(N,\mathbf{G};\eta):=\frac{2}{N}\sum_{t=N/2}^{N-1}\mathbf{w}_{t}.$ (3.2)
@@
General set-up. In this study, we compare the generalization performance of SGD and ridge regression from the standpoint of excessive risk for the same amount of training samples, which essentially is the notion of Bahadur statistical efficiency (Bahadur, 1967). Specifically, our goal is to demonstrate the existence of a preconditioning matrix (through construction) that enables SGD to achieve a comparable or superior excessive risk to ridge regression when provided with an equal amount of training data. To ensure a meaningful comparison, we focus on a setting where both SGD and ridge regression are considered “generalizable”, meaning they achieve a vanishing excess risk when $n\rightarrow\infty$ (note that $d$ can also approach $\infty$) with the optimal hyperparameter configurations. For this purpose, in the forthcoming sections, we assume sufficient training samples (a large enough $N$) and a constant level signal-to-noise ratio, i.e., $\rho^{2}=\frac{\|\mathbf{w}^{*}\|^{2}_{\mathbf{H}}}{\sigma^{2}}=\Theta(1).$

### v3

@@ Abstract
Stochastic gradient descent (SGD) exhibits strong algorithmic regularization effects in practice and plays an important role in the generalization of modern machine learning. However, prior research has revealed instances where the generalization performance of SGD is worse than ridge regression due to uneven optimization along different dimensions. Preconditioning offers a natural solution to this issue by rebalancing optimization across different directions. Yet, the extent to which preconditioning can enhance the generalization performance of SGD and whether it can bridge the existing gap with ridge regression remains uncertain. In this paper, we study the generalization performance of SGD with preconditioning for the least squared problem. We make a comprehensive comparison between preconditioned SGD and (standard & preconditioned) ridge regression. Our study makes several key contributions toward understanding and improving SGD with preconditioning. First, we establish excess risk bounds (generalization performance) for preconditioned SGD and ridge regression under an arbitrary preconditions matrix. Second, leveraging the excessive risk characterization of preconditioned SGD and ridge regression, we show that (through construction) there exists a simple preconditioned matrix that can make SGD comparable to (standard & preconditioned) ridge regression. Finally, we show that our proposed preconditioning matrix is straightforward enough to allow robust estimation from finite samples while maintaining a theoretical improvement. Our empirical results align with our theoretical findings, collectively showcasing the enhanced regularization effect of preconditioned SGD. footnotetext: $\dagger$: corresponding authors
@@ 1 Introduction
A growing body of work has emerged, dedicated to understanding and characterizing the generalization performance of SGD in the context of the linear regression problem. These studies have unveiled a close relationship between the implicit regularization inherent to SGD and the explicit $L_{2}$-type regularization (ridge) (Zhang et al., 2017; Tikhonov, 1963; Ali et al., 2019, 2020; Gunasekar et al., 2018; Suggala et al., 2018; Zou et al., 2021a). In particular,  (Zou et al., 2021a) has presented an instance-based comparison study between the generalization performance (measured by excessive risk) of SGD and ridge regression. They show that (single-pass) SGD can achieve comparable generalization performance than that of ridge regression in certain problem instances but fall behind in others.111we provide concrete examples for further illustration in Appendix F.
A notable insight derived from the comparative analysis conducted by (Zou et al., 2021a) is that the underperformance of SGD in certain scenarios can be attributed to an inherent imbalance optimization within the data covariance matrix. This imbalance results in a substantial bias error, primarily associated with the “relatively small” eigenvalues of the data covariance matrix. To address this limitation of SGD, preconditioning emerges as a natural solution. Preconditioning is a commonly used technique to enhance the performance of the optimizer (Benzi, 2002; Avron et al., 2017; Gupta et al., 2018; Amari et al., 2020; Kang et al., 2023; Li, 2017; Li et al., 2016); it involves modifying the update steps of the algorithm by transforming the data space, offering a mechanism to rebalance the optimization process across different directions. Despite extensive research that has explored the implicit regularization of SGD and its parallels with ridge regression, the realm of preconditioned SGD remains relatively uncharted. Recognizing this gap in the current body of knowledge, our study investigates the implicit regularization of SGD when augmented with preconditioning. We adhere to the standard settings (least square problem) used in the existing literature (Zou et al., 2021a, b, 2023; Tsigler & Bartlett, 2023) and aim to address the following central question of our study:
@@
In essence, our objective is to investigate whether a thoughtfully designed preconditioning matrix can empower SGD to be consistently comparable with ridge regression in the least square problem. Furthermore, we aim to delve deeper into the dynamic relationship between SGD and ridge regression when both methods incorporate preconditioning. Achieving these goals presents several key challenges. First and foremost, existing results characterizing the excessive risk of SGD and ridge regression predominantly pertain to scenarios without preconditioning. The introduction of preconditioning alters the learning dynamics of SGD and ridge regression, potentially leading to distinct characterizations of their excessive risk. Secondly, as discussed earlier, the effectiveness of preconditioning for SGD is contingent upon its impact on the eigenspectrum of the data covariance matrix. A naive design of the preconditioning matrix may fail to align with the excessive risk characteristics of SGD, potentially leading to adverse consequences. Lastly, obtaining precise information regarding the eigenspectrum of the covariance matrix can be impractical in real-world settings. To ensure the practical applicability and validity of the preconditioning technique in such scenarios, we must consider a design that remains sufficiently straightforward to facilitate robust estimation.
@@
In this paper, we show (through construction) that there exists a preconditioning design that allows the excessive risk of SGD to be consistently comparable with ridge regression. In addition, we establish corresponding analytical tools to tackle the challenges stated earlier. The main contributions of our work are summarized as follows.
@@
  • Building upon the insights derived from the analysis of excessive risk, we propose a simple (yet effective) preconditioning matrix tailored to SGD, which leverages information from the data covariance matrix. We show that under the theoretical setting (with access to the precise information of data covariance matrix), the excessive risk of SGD with our proposed preconditioning consistently matches that of standard ridge regression and a representative family of preconditioned ridge regression (Theorems 4.3 &  4.4). This shows that preconditioning can bridge the existing gap between SGD and ridge regression, supporting the assertion of the central question above.
  • In practical scenarios where precise information about the data covariance matrix is unavailable, our study illustrates that our proposed preconditioning design allows for robust estimation using a finite set of inexpensive unlabeled data. Remarkably, even when we employ the estimated version of our proposed preconditioning matrix, the excessive risk of SGD can still consistently match that of (standard & preconditioned) ridge regression (Theorems 4.5 &  4.6). This discovery not only highlights the significance of SGD in practical machine learning applications but also underscores the practical advantages offered by preconditioning.
Notation. We use lowercase letters to denote scalars, and boldface letters to denote vectors and matrices, respectively. For a vector $\mathbf{x}$, $\mathbf{x}[i]$ denotes the $i$-th coordinate of $\mathbf{x}$. For two functions $f(x)\geq 0$ and $g(x)\geq 0$ defined on $x>0$, we write $f(x)\lesssim g(x)$ if $f(x)\leq c\cdot g(x)$ for some absolute constant $c>0$; we write $f(x)\gtrsim g(x)$ if $g(x)\lesssim f(x)$; we write $f(x)\eqsim g(x)$ if $f(x)\lesssim g(x)\lesssim f(x)$. For a vector $\mathbf{w}\in\mathbb{R}^{d}$ and a positive semidefinite matrix $\mathbf{H}\in\mathbb{R}^{d\times d}$, we denote $\|\mathbf{w}\|^{2}_{\mathbf{H}}:=\mathbf{w}^{\top}\mathbf{H}\mathbf{w}$.
@@ 2 Related Work
We provide an overview of the technical advancements in the analysis of ridge regression and SGD, as well as related research comparing SGD to explicit norm-based regularization and ridge regression. Additionally, we discuss the current usage of preconditioning in both SGD and ridge regression.
@@
Excessive Risk Bounds of Ridge Regression. Ridge regression holds a central position in the field of machine learning (Hsu et al., 2012; Kobak et al., 2020; Tsigler & Bartlett, 2023; Hastie et al., 2009). In the underparameterized scenario, there exists comprehensive understanding on the excess risk bounds of ridge regression (Hsu et al., 2012). However, in the overparameterized regime, a large body of work focuses on characterizing the excess risk of ridge regression in the asymptotic scenario where both the sample size and the dimension approach infinity with a constant ratio (Dobriban & Wager, 2018; Hastie et al., 2022; Wu & Xu, 2020; Xu & Hsu, 2019). Recent advancements by (Tsigler & Bartlett, 2023) have provided detailed non-asymptotic risk bounds for ordinary least squares and ridge regression in overparameterized contexts. Notably, these bounds depend on the data covariance spectrum and are applicable even when the ridge parameter is zero or negative. This body of work forms a foundation of the theoretical framework for our paper.
@@
Excessive Risk Bounds of SGD. In the finite-dimensional setting, risk bounds for one-pass, constant-stepsize SGD have been well-established (Dieuleveut et al., 2017; Bach & Moulines, 2013; Jain et al., 2017, 2018; Défossez & Bach, 2015; Paquette et al., 2022). In the under-parameterized setting, constant-stepsize SGD with tail-averaging has been recognized for achieving the minimax optimal rate for least squares (Jain et al., 2017, 2018). A recent study by (Zou et al., 2023) expanded on the previous analysis, offering a detailed risk bound for SGD in the overparameterized regime. This work provides closely aligned upper and lower excess risk bounds for constant-stepsize SGD, distinctly defined in relation to the complete eigenspectrum of the population covariance matrix. These findings are critically influential in the context of our paper.
@@
Implicit Regularization of SGD and Explicit Regularization. It is well known that in the least squares problems, multi-pass SGD approaches the solution with the minimum norm, representing a recognized implicit bias of SGD (Neyshabur et al., 2014; Zhang et al., 2017; Gunasekar et al., 2018). However, this norm-based regularization alone does not fully explain the optimization behaviour of SGD in more extensive scenarios such as convex but non-linear models (Arora et al., 2019; Dauber et al., 2020; Razin & Cohen, 2020). Recent work by (Zou et al., 2021a) presents a comparative analysis between SGD and ridge regression from the perspective of excessive risk, demonstrating that SGD can match the performance of ridge regression in some scenarios but fall short in others. Their result has motivated us to explore SGD with preconditioning. In this paper, we advance the understanding of SGD by exploring its implicit regularization in the presence of preconditioning.
@@
Preconditioning in SGD and ridge. Preconditioning is a widely employed optimization technique (Benzi, 2002) for enhancing the efficiency and effectiveness of optimization algorithms. Its primary objective is to transform complex or ill-conditioned optimization problems into more manageable forms. Preconditioning techniques have been extensively explored in both SGD and ridge regression, serving various purposes (Benzi, 2002; Li, 2017; Gonen et al., 2016; Amari et al., 2020; Xu et al., 2023; Kang et al., 2023). However, the impact of preconditioning on the excessive risk of SGD and ridge regression, and whether it can enable SGD to consistently match the performance of ridge regression, remain open questions that motivate our research.
@@ 3 Problem Setup and Preliminaries
In this paper, our objective is to close an open question in the existing learning literature. To do so, we explore the usage of precondition in SGD, adhere to the setting in previous studies, and compare the generalization performance of preconditioned SGD and ridge regression in the context of solving least square problems in the overparameterized regime 222we present a further discussion on our setting and how our results can be extended in Appendix E. We denote a feature vector in a separable Hilbert space $\mathcal{H}$ as $\mathbf{x}\in\mathcal{H}$. The dimensionality of $\mathcal{H}$ is denoted as $d$, which can be infinite-dimensional when applicable. We use $y\in\mathbb{R}$ to denote a response and it is generated as follows:
@@
where $\mathbf{w}^{*}\in\mathcal{H}$ represents the unknown true model parameter, and $\epsilon\in\mathbb{R}$ is the model noise. The goal of the least squares problem is to estimate the true parameter $\mathbf{w}^{*}$ by minimizing the following:
$L(\mathbf{w}^{*})=\min_{\mathbf{w}\in\mathcal{H}}L(\mathbf{w}),\quad\text{where}\quad L(\mathbf{w}):=\frac{1}{2}\mathbb{E}_{(\mathbf{x},y)\sim\mathbb{D}}\big{[}(y-\langle\mathbf{w},\mathbf{x}\rangle)^{2}\big{]},$
where $\mathbb{D}$ represents the data distribution. The generalization performance of an estimated $\mathbf{w}$ found by an algorithm, whether it is SGD or ridge regression, is evaluated based on the excessive risk:
@@
There exists a positive constant $\alpha$ such that for any PSD matrix $\mathbf{A}$, it holds that: $\mathbb{E}[\mathbf{x}\mathbf{x}^{\top}\mathbf{A}\mathbf{x}\mathbf{x}^{\top}]-\mathbf{H}\mathbf{A}\mathbf{H}\preceq\alpha\mathop{\mathrm{tr}}(\mathbf{A}\mathbf{H})\cdot\mathbf{H}.$
@@
There exists a positive constant $\sigma$ such that: $\mathbb{E}_{\mathbf{x},\epsilon}[\epsilon^{2}\mathbf{x}\mathbf{x}^{\top}]\leq\sigma^{2}\mathbf{H}.$
To analyze $\mathbf{w}^{*}$ and $\mathbf{H}$ in terms of excessive risk, we introduce the following notation:
@@
Preconditioned SGD. We consider single-pass preconditioned SGD with a constant step size and tail-averaging (Bach & Moulines, 2013; Jain et al., 2017, 2018). At the $t$-th iteration, a fresh example $(\mathbf{x}_{t},y_{t})$ is sampled from the data distribution and to perform the update:
${\mathbf{w}}_{t+1}=\mathbf{w}_{t}-\eta\cdot{\color[rgb]{0,0,1}{\mathbf{G}}}\nabla l(\mathbf{w}_{t};\mathbf{x}_{t},y_{t})={\mathbf{w}}_{t}-\eta\cdot(\langle{\mathbf{w}}_{t},\mathbf{x}_{t}\rangle-y_{t})\cdot{\color[rgb]{0,0,1}{\mathbf{G}}}\mathbf{x}_{t}$ (3.1)
where $\eta>0$ is a constant stepsize (also referred to as the learning rate) and $\mathbf{G}$ (highlighted in blue) is the preconditioned matrix — the key difference from the standard SGD. Note that if $\mathbf{G}=\mathbf{I}$, then the update rule (3.1) reduces to the standard SGD with constant step size.
@@
$\mathbf{w}_{\mathrm{sgd}}(N,\mathbf{G};\eta):=\frac{2}{N}\textstyle{\sum_{t=N/2}^{N-1}}\mathbf{w}_{t}.$ (3.2)
@@
General set-up. In this study, we compare the generalization performance of SGD and ridge regression from the standpoint of excessive risk for the same amount of training samples, which essentially is the notion of Bahadur statistical efficiency (Bahadur, 1967). Specifically, our goal is to demonstrate the existence of a preconditioning matrix (through construction) that enables SGD to achieve a comparable or superior excessive risk to ridge regression when provided with an equal amount of training data. For comparison, we adhere to the previous studies and adopt the order-based comparison framework (which focuses on growth rate and is a common framework used in similar theoretical studies). To ensure a meaningful comparison, we focus on a setting where both SGD and ridge regression are considered “generalizable”, meaning they achieve a vanishing excess risk when $n\rightarrow\infty$ (note that $d$ can also approach $\infty$) with the optimal hyperparameter configurations. For this purpose, in the forthcoming sections, we assume sufficient training samples (a large enough $N$) and a constant level signal-to-noise ratio, i.e., $\rho^{2}=\frac{\|\mathbf{w}^{*}\|^{2}_{\mathbf{H}}}{\sigma^{2}}=\Theta(1).$